In the decomposition of H2O2 at 300K, the energy of activation was found to be 18 kcal/mol while it decreases to 6kcal/mol when the decomposition was carried out in the presence of a catalyst at 300k. How many times is the catalysed reaction faster than the uncatalyased one?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

K = Ae-E/RT                  (E is activation energy)

 

rate(R)= K(a)n                             (k=rate constant & a is concentration , n is order of reaction)

 

rate directily proportional to rate constant so

 

R1/R2 = K1/K2 = e[-E1+E2]/RT

R1/R2 = e[-18+6]/RT                           (R=.08321units , T=300)

          =e-[0.5]                               (e=2.8 )

R2 = 1.67R1               (approx)